Problem 3 A parallel-plate capacitor with plate area 5.00 cm2 and air-gap separation 0.29 mm is connected to a 12.00 V battery, and fully charged. The battery is then disconnected (a) What is the charge on the capacitor? 1.83x10-10 (c You are correct. Computers answer now shown above Your receipt is 160-9143 Previous Tries (b) The plates are now pulled to a separation of 0.48 mm. What is the charge on the capacitor now? 183x1010 (c You are correct. Computers answer now shown above Previous Tries Your receipt is 160-7735 (c) What is the potential difference across the plates now? 1.99×101 V You are correct. Computers answer now shown above Your receipt is 160-6675 . Previous Tries (d) How much work was required to pull the plates to their new separation?
Answer #1

Solution) (d) work done W = ?

Charge q = 1.83×10^(-10) C =Q1

W = (1/2)q(dV)

dV is change in potential

We have Q =CV

Q1 = C1V1

C1 = Q1/V1 = (1.83×10^(-10))/(12)

C1 =0.152×10^(-10) F

C = (epsilonnot)A/d

C is inversely proportional to d

New capacitance with new separation

C2 = C1(d1/d2)

C2 = (0.152×10^(-10))(0.29/0.48)

C2=0.091×10^(-10) F

V2 = Q2/C2 = (1.83×10^(-10))/(0.091×10^(-10))

V2 = 20.10 V

dV = V2 - V1 = 20.10 - 12 = 8.1 V

W = (1/2)(1.83×10^(-10))(8.1)

W=7.41×10^(-10) J